OpenStudy (anonymous):

let f(x)=│x-3│ which of the following conditional statement hold true A.f(x)=-(x-3) B.f(x)0 if x=0 C.f(x)=x-3 if x>0 D.f(x)=x-3 if x>3

OpenStudy (anonymous):

@satellite73 can u help

OpenStudy (sidsiddhartha):

if f(x)=|x| then f(x)=x ,when x>0 f(x)=-x,when x<0 use this property

OpenStudy (anonymous):

can u help me solve it

OpenStudy (unklerhaukus):

consider x is 5 is A correct?

OpenStudy (anonymous):

yea

OpenStudy (unklerhaukus):

check that f(x)=│x-3│ f(5)=│5-3│ = ? A. f(x)=-(x-3) f(5)=-(5-3) = ?? are ? and ?? equal

OpenStudy (anonymous):

ohh

OpenStudy (unklerhaukus):

what do you get for ? and ??

OpenStudy (anonymous):

im trying to figure it out

OpenStudy (anonymous):

can u help me

OpenStudy (unklerhaukus):

f(5)=│5-3│ what is 5-3

OpenStudy (anonymous):

2

OpenStudy (unklerhaukus):

f(5)=│5-3│ = | 2| what is absolute value of 2

OpenStudy (anonymous):

2

OpenStudy (unklerhaukus):

good so f(x)=│x-3│ f(5)=│5-3│ = | 2 | = 2 A. f(x) = -(x-3) f(5) = -(5-3) = -(2) are 2 and -(2) equal? could A. be correct?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

good, so we can rule out option A.

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

lets look at option B. now f(x)=│x-3│ B. f(x) = 0 if x=0 lets compute f(0) f(0) =│0-3│ = what does this simplify to become?

OpenStudy (anonymous):

3

OpenStudy (unklerhaukus):

thats right f(0) = 3 is this that same as B. f(x) = 0 if x=0 ?

OpenStudy (anonymous):

yea

OpenStudy (unklerhaukus):

3 is the same as 0?

OpenStudy (anonymous):

oh noo

OpenStudy (unklerhaukus):

3 and 0 are not the same are they, so can option B. be the correct one?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

f(x)=│x-3│ How about option C. C. f(x) = x-3 if x>0 lets try x = 7 because 7>0 what is f(7) =

OpenStudy (anonymous):

7

OpenStudy (unklerhaukus):

f(7) =│7-3│ = . . .

OpenStudy (anonymous):

4

OpenStudy (unklerhaukus):

ok and C. f(7) = 7-3 = . . .

OpenStudy (anonymous):

4

OpenStudy (unklerhaukus):

that's right so we tested x=7 and found no problem with C. but what if we tested x = 2 because 2 is still greater than zero f(x) =│x-3│ f(2) = | 2-3 | = C. f(x) = x - 3 f(2) = 2 - 3 = are these the same?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

what did you get for f(2) = | 2-3 | =

OpenStudy (anonymous):

1

OpenStudy (unklerhaukus):

yeah, and what about C. f(x) = x - 3 f(2) = 2 - 3 =

OpenStudy (anonymous):

so c is correct

OpenStudy (unklerhaukus):

what did you get for 2 - 3

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so it has to Be D

OpenStudy (unklerhaukus):

is 1 the same as -1 ?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

so now lets check D. f(x)=│x-3│ D. f(x)=x-3 if x>3 what is f(4) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4-3=1

OpenStudy (anonymous):

│4-3│=1

OpenStudy (unklerhaukus):

that fits with f(x)=│x-3│ and with D. f(x)=x-3 if x>3 what about f(23)?

OpenStudy (anonymous):

│23-3│=20

OpenStudy (unklerhaukus):

and D. f(x) = 23-3 (23>3) = 20 also

OpenStudy (anonymous):

so its D

OpenStudy (unklerhaukus):

yep \[\checkmark\]

OpenStudy (anonymous):

can u help me with1 more

OpenStudy (unklerhaukus):

i think i have to go to bed, soory

OpenStudy (anonymous):

it wont take long plx i just beed anwser

OpenStudy (unklerhaukus):

go on then

OpenStudy (anonymous):

let f(x)= -2x x<0 0 x>0 which is equivalent A.f(x)=x+│x│ B.f(x)=│x│-x C.f(x)=2x D.f(x)=2│x│

OpenStudy (unklerhaukus):

test some numbers for x

OpenStudy (anonymous):

what do u think the anwser is

OpenStudy (unklerhaukus):

what if x=1 ? (1>0) is A. correct?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

wait no

OpenStudy (unklerhaukus):

can you show some working

OpenStudy (anonymous):

i only have 5 min to finish and this is the last question

OpenStudy (unklerhaukus):

im not going to tell you the answer, but i can check your working

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got B is that corect

OpenStudy (unklerhaukus):

where is the working?